Find the equation of a straight line passing through points P(1,3) and Q(-2,0) and express it in parametric form.

Grade:6

1 Answers

Shivanshu Amitabh
6 years ago
Here, x1= 1 , x2=-2 ,y1=3 , y2=0.Usinv two point form, y-y1 = y2-y1/x2-x1(x-x1)Putting the values,y-3 = 0-3/-2-1(x-1)y-3=-3/-3(x-1)y-3=x-1y=x-1+3[y=x+2] is the equation of the given line.To convert it into parametric form, let y=f(x)Now, let x=t, theny=f(t)y=t+2Therefore, x=t and y=t+2 is the parametric form of given line.